Database last updated on Thu, 29 Jan 2026 05:40:53 +0000 by localhost

Critical warnings from b-files' comments for days between 2024337 and 2024337 (1 sec, 6 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2024 337 05 06 27 New value set for watermark SUPPLY.VOLTAGE.HIGH123B33724.254TORON#25405:09:541-1
rl:2024 337 05 30 01 New value set for watermark LOOP.RUN.TIME: 680.999994 millisec.83B33724.254TORON#25405:09:591-1
rl:2024 337 06 55 36 PMT counter failure11B33724.254TORON#25405:10:181-1
hp:Spectrometer alignment trouble.21B33724.254TORON#25411:26:081-1
hg:Lamp not on... hg test terminated F1= 0 F4= 2011B33724.254TORON#25422:21:011-1
hg:Lamp not detected f1= 0 f4= 2011B33724.254TORON#25422:21:021-1

Instrument constant changes for the week before 2024337 (4 sec, 0 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
254202433741020.0

HP with step corrections larger than 20 steps (0 sec, 1 rows)

SerialYYYYJJJHP largemax correction
25420243371279.9

Non-critical warnings from b-files' comments for days between 2024337 and 2024337 (0 sec, 4 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2024 337 05 05 29 Reset Warm -> Tepid from 2024 337 05 05 23 until 2024 337 05 05 23 . Counts: 1,023B33724.254TORON#25405:09:511-1
dp:Doing dispersion with internal HG.199B33724.254TORON#25411:44:391-1
skc:HOME key pressed299B33724.254TORON#25413:04:101-1
pd:Going to menu199B33724.254TORON#25413:13:441-1

Measurements and tests summary.


_jday_loc_serial_skcdtedhghprssluv
337TORONTO#254HOME key pressed,inside,menu,Must do HG now,Waiting until ZA = 170 now 128.7,Waiting until ZA = 170 now 130.5,Waiting until ZA = 170 now 132.3,Waiting until ZA = 170 now 134.1,Waiting until ZA = 170 now 135.9,Waiting until ZA = 170 now 137.6,Waiting until ZA = 170 now 139.3,Waiting until ZA = 170 now 141.1,Waiting until ZA = 170 now 142.8,Waiting until ZA = 170 now 144.4,Waiting until ZA = 170 now 146,Waiting until ZA = 170 now 147.6,Waiting until ZA = 170 now 149.1,Waiting until ZA = 170 now 150.5,Waiting until ZA = 170 now 151.9,Waiting until ZA = 170 now 153.2,Waiting until ZA = 170 now 154.3,Waiting until ZA = 170 now 155.4,Waiting until ZA = 170 now 156.3,Waiting until ZA = 170 now 157.1,Waiting until ZA = 170 now 157.7,Waiting until ZA = 170 now 158.1,Waiting until ZA = 170 now 158.3,Waiting until ZA = -110 now -111.6,Waiting until ZA = -110 now -113.4,Waiting until ZA = -110 now -115.2,Waiting until ZA = -110 now -117,Waiting until ZA = -110 now -118.8,Waiting until ZA = -110 now -120.6,Waiting until ZA = -110 now -122.5,Waiting until ZA = -110 now -124.3,Waiting until ZA = -110 now -126.1,Waiting until ZA = -110 now -127.9,Waiting until ZA = -110 now -129.7,Waiting until ZA = -110 now -131.5,Waiting until ZA = -110 now -133.3,Waiting until ZA = -110 now -135.1,Waiting until ZA = -110 now -136.9,Waiting until ZA = -110 now -138.6,Waiting until ZA = -110 now -140.3,Waiting until ZA = -85 now -86.1,Waiting until ZA = -85 now -87.5,Waiting until ZA = -85 now -88.9112305111111
(0 sec)

Number of days since last dome/window cleaning (0 sec, 1 rows)

Missing files (0 sec, 1 rows)

Seriallast datadays ago
25420243380-1

Questionable Run/Stop results. (0 sec, 1 rows)

SerialJdayslit 0dark rsdark countsslit 1slit 2slit 3slit 4slit 5slits 3+5
25420243371.0002.0002 / 0 (0.6M)0.9971.0000.9991.0001.0000.999

Run/Stop results for last week. (0 sec, 1 rows)

SerialTotalFailed 0Failed darkFailed 1Failed 2Failed 3Failed 4Failed 5Failed double
25462010100000

AP results for SB Brewers. (0 sec, 1 rows)

SerialJdayt1 offt1 ont2 offt2 ont3 offt3 onHV offHV on+12V off+12V on+5V off+5V on-12V off-12V on+24V off+24V on+5V sps off+5V sps onSL A offSL A onSL V offSL V onWatts
254202433728.328.429.229.027.427.31388.81388.812.112.15.15.1-12.0-12.024.324.35.05.00.01.50.010.2015.8

Suspicious HG test results (correlation<0.98 or counts<5000). (0 sec, 1 rows)

SerialJdaycountPeakmin.corr.max.corr.30-day history
#2543375710400